80 likes | 154 Views
Detailed overview of progress and challenges faced during the TOF modules installation at Inti Lehmann University of Glasgow in May 2011. Includes realignment observations, solutions for misalignments, and pending tasks to be completed.
E N D
Status of the TOFs Inti Lehmann University of Glasgow 27 June 2011
Work done in May • All scintillators • detached • re-aligned • secured with new brackets • Short flasher fibres attached • Modules moved to Doris • Framing done • Subframes moved into main frame • All!? 34 TOF modules installed OLYMPUS TOF Status 27/06/2011
May2011 OLYMPUS TOF Status 27/06/2011
Misalignment realised • Observed effect • Chambers > TOF 17 • Gap from target • Huge mismatch: 10-25cm • TOF CAD • On-site (Michael) • Explanation • Unsynchronised changes in CAD • Solution • Move group 3 (modifying brackets, add. support) • Add TOFs #18 (add. cables, short by 1 in flasher) OLYMPUS TOF Status 27/06/2011
Not beautiful but... OLYMPUS TOF Status 27/06/2011
Status and to-do • 34 TOFs in place • 2 TOFs in preparation • Hardware to extend frames expected on Thu • 72 HV cables and 70 signal cables labelled • Require ordering more signal cables • Route cables to hut • 72 mu-metal shields ready • Mounting on top may prove difficult • 7 (possibly) loose PMTs • Re-attachment started • Flasher work • Light-coupler, electronics, route • Tests OLYMPUS TOF Status 27/06/2011
Mismatch OLYMPUS TOF Status 27/06/2011